Using traditional methods it takes 106.0 hours to receive a basic driving license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique on 60 students and observed that they had a mean of 107.0 hours. Assume the variance is known to be 49.00. Is there evidence at the 0.02 level that the technique performs differently than the traditional method?
Answer #1

To Test :-
H0 :- µ = 106
H1 :- µ ≠ 106

Test Statistic :-
Z = ( X̅ - µ ) / ( σ / √(n))
Z = ( 107 - 106 ) / ( 7 / √( 60 ))
Z = 1.1066


Test Criteria :-
Reject null hypothesis if | Z | > Z( α/2 )
Critical value Z(α/2) = Z( 0.02 /2 ) = 2.326 ( From Z table )
| Z | > Z( α/2 ) = 1.1066 < 2.326
Result :- Fail to reject null hypothesis

Decision based on P value
Reject null hypothesis if P value < α = 0.02 level of significance
P value = 2 * P ( Z > 1.1066 ) = 2 * 1 - P ( Z < 1.1066 ) = 0.2685
Since 0.2685 > 0.02 ,hence we fail to reject null hypothesis
Result :- We fail to reject null hypothesis

There is insufficient evidence to support the claim that new technique performs differently than the traditional method.